P(-3.10<z<1 90)

1
Expert's answer
2022-02-07T17:18:11-0500

"P(-3.10<Z<1.90)=P(Z<1.90)-P(Z<-3.10)="

"=P(Z<1.90)+P(Z<3.10)-1=0.9703."